OpenStudy (matt6288):

If a sample of Mg(NO3)2 has a mass of 160.55 grams, how many formula units of Mg(NO3)2 are present in this sample?

OpenStudy (matt6288):

is this written correctly?

OpenStudy (matt6288):

160.55g x Mg(NO3)2/148.33g x 6.022x10^23 Formula units/Mg(NO3)2 = 6.5181 x 10^23 Formula units

OpenStudy (matt6288):

That is exactly what I thought it was... than I saw the answer and it was, "6.5181 x 10^23 Formula units" no work just answer.

OpenStudy (aaronq):

"formula unit" is the name used for inorganic compounds represented as their chemical formulas are written (e.g. \(Mg(NO_3)_2\), NaCl), which as reductions to a lowest common factor (it's not \(Na_2Cl_2\)) as they exist in large clusters, or networks, and not as single molecules. https://www.wikiwand.com/en/Formula_unit So for this question you would find the moles of Mg(NO3)2 and multiply by avogadro's number. That gives you the number of formula units. \( \sf moles=\dfrac{mass}{Molar~mass}\) \(\sf moles=\dfrac{particles}{Avogadro′s number}\) for this question formula units would represent particles.